Using the energy conservation law determine the following...
If the minimum velocity needed by the skater to pass the loop is 5.42m/s, what is the minimum initial height that allows the skater complete the loop?
1
Expert's answer
2020-04-22T10:05:04-0400

So that skater does not leave the loop it is necessery that "w_n=v^2/R\geq g" then "v_{min}=\sqrt{gR}"

and "R=v^2_{min}\/g." As "v^2_{min}/2+mg\sdot2R=mgH" hence "v^2_{min}\/2=g(H-2v^2_{min}\/g)"

hence "H=\frac{5v^2_{min}}{2g}" then "H\approx7.5m"